OpenStudy (anonymous):

If r(t)=cos(−3t)i+sin(−3t)j+7tk, compute the tangential and normal components of the acceleration vector.

OpenStudy (amistre64):

is r(t) the position vector?

OpenStudy (amistre64):

r'(t) = velocity = <3sin(t),-3cos(t),7> r''(t) = <3cos(t),3sin(t),0> maybe

OpenStudy (amistre64):

im sure that simplictic and i forgot a few things

OpenStudy (anonymous):

Yeah, r(t) is the position vector. But is the tangential just the first derivative and the normal is just the second derivative?

OpenStudy (amistre64):

well, i am sure the first is the tangent vector; im shaky as to the 2nd tho

OpenStudy (amistre64):

i dont see why not, but still its been too long to remember it

OpenStudy (anonymous):

I'll try and let you know. Thanks!

OpenStudy (anonymous):

It says it needs a formula that returns a number. But what we gave it returns a list of numbers.

OpenStudy (amistre64):

let me read up on it and see what i can recall

OpenStudy (amistre64):

you have any idea in what direction we can go with it?

OpenStudy (amistre64):

maybe like finding curvature and whatnot?

OpenStudy (anonymous):

I mean, I don't feel like it would be curvature, but maybe it is. I'm thinking maybe do a magnitude of the vectors you just gave me?

OpenStudy (amistre64):

this is what im thinking of... as from http://omega.albany.edu:8008/calc3/curves3-dir/lecture.html v = dR/dt = dR/ds * ds/dt = T * ds/dt a = dv/dt = dT/dt * ds/dt + T * = = dT/dt * (ds/dt)^2 + T * dT/ds = K*N, where K is the radius of curvature. a = K(ds/dt)^2 * N + * T a = aN * N + aT * T aN = K*(ds/dt)^2 = K |v|^2 = (|v|^2)/rho, where rho is the radius. aT = = d|v|/dt aN from above is the centripetal acceleration.

OpenStudy (amistre64):

its prolly more readable from this tho: http://www.ltcconline.net/greenl/courses/202/vectorFunctions/tannorm.htm

OpenStudy (anonymous):

checking it out now. thanks

OpenStudy (amistre64):

\begin{array}l r'(t) = velocity = <3sin(t),-3cos(t),7>\\\\ |r'(t)|=\sqrt{9sin^2(t)+9cos^2(t)+49}\\\\ |r'(t)|=\sqrt{9+49}\\\\ |r'(t)|=\sqrt{58}\\\\ \\\\ r''(t) = <3cos(t),3sin(t),0>\\\\ |r''(t)| = \sqrt{9cos^2(t)+9sin^2(t)}\\\\ |r''(t)| = \sqrt{9}=3\\\\ \end{array}

OpenStudy (anonymous):

Neither of those just worked :(. I just tried this for the first part: (cos(-3t)+(sin(-3t))+(7t))/(sqrt(((3sin(t))+(-3cos(t))+(7))))

OpenStudy (amistre64):

http://tutorial.math.lamar.edu/Classes/CalcIII/Velocity_Acceleration.aspx here is another site that helps me out

OpenStudy (amistre64):

\[a=a_tT+a_nN\] where T and N are the unit vectors

OpenStudy (amistre64):

\[a_t = \frac{r'*r''}{|r'|}\] \[a_t = \frac{<3sin(t),-3cos(t),7>*<3cos(t),3sin(t),0>}{|\sqrt{58}|}\] \[a_t = \frac{9sin(t)cos(t)-9sin(t)cos(t)+0}{\sqrt{58}}\] \[a_t = \frac{0}{\sqrt{58}}=0\] but I gotta wonder

OpenStudy (anonymous):

you were right!!

OpenStudy (anonymous):

That's the first one. Awesome, thank you!

OpenStudy (amistre64):

\[a_n=\frac{|r'x\ r''|}{|r'|}\] \[a_n=\frac{|<3sin(t),-3cos(t),7>x<3cos(t),3sin(t),0>|}{|\sqrt{58}|}\] \[a_n=\frac{|-21sin(t)+21cos(t)+9|}{\sqrt{58}}\] \[a_n=\frac{|21[cos(t)-sin(t)]+9|}{\sqrt{58}}\] if i did the cross right, this is what i get for the normal component

OpenStudy (amistre64):

might have to put some ijks in it tho

OpenStudy (amistre64):

\[a_n=\frac{|-21sin(t)\ I+21cos(t)J+9K|}{\sqrt{58}}\]

OpenStudy (amistre64):

ugh, i think the top there is a cross of magnitudes

OpenStudy (anonymous):

yeah, I'm trying different things and what you put up before but it hasn't worked so far.

OpenStudy (amistre64):

\[a_n=\frac{\sqrt{(-21sin(t))^2\ +(21cos(t))^2+9^2}}{\sqrt{58}}\] simplified

OpenStudy (amistre64):

sin^2+cos^2 = 1 so we are left up top with: \[\sqrt{21^2+9^2}\]

OpenStudy (amistre64):

forgot yo add 81 up top

OpenStudy (amistre64):

\[a_n=\frac{\sqrt{522}}{\sqrt{58}}\]

OpenStudy (amistre64):

which is 3 :)

OpenStudy (anonymous):

:(...none of those last 5 answers have worked so far.

OpenStudy (amistre64):

bummer, cause unless i read it wrong, thats the best I got for it

OpenStudy (anonymous):

you've done plenty! I'll keep going at it alone. Thanks for all your help!

OpenStudy (amistre64):

good luck with it ;)